The speed of a boat in still water is 16 mph. The boat traveled the same distance downstream in 9 minutes that it traveled upstream in 15 minutes. Find the speed of the current.
Mary

Answer by Paul(988)

Let the speed of the current be x
9 minutes = 0.15hours
15 minutes = 0.25 hors
EquatioN:

0.15(16+x)=0.25(15-x)
2.4+0.15x=4-0.25x
-1.6=-0.4x
x=4

Hence, the speed of the current is 4mph.
Paul.